### v2

@@ Abstract
In this note, we consider infinite horizon optimal control problems with deterministic systems. Since exact solutions to these problems are often intractable, we propose a parallel model predictive control (MPC) method that provides an approximate solution. Our method computes multiple lookahead minimization problems at each time, where each minimization may involve a different number of lookahead steps, and terminal cost and constraint. The parallel MPC method then applies the first control of the lookahead minimization with the lowest cost. We show that the proposed method has better performance guarantee than the MPC scheme involving a single lookahead minimization. Moreover, we establish the validity of a parallel implementation of our method in the presence of multiple computing units.
@@ 1 INTRODUCTION
$\mu^{*}(x)\in\arg\min_{u\in U(x)}\big{\{}g(x,u)+J^{*}\big{(}f(x,u)\big{)}\big{\}},$ (3)
if the minimum in (3) can be attained for all $x$.
In the context of dynamic programming (DP), the goal is to obtain an optimal policy $\mu^{*}$ and optimal cost function $J^{*}$. However, computing $\mu^{*}$ and $J^{*}$ is often intractable. To obtain an approximate solution, one may use some known function $\bar{J}$ in place of $J^{*}$ and perform a minimization calculation similar to (3) at each time. If $\bar{J}$ resembles $J^{*}$ in some way, we may expect the obtained policy $\tilde{\mu}$ close to optimal, in the sense that $J_{\tilde{\mu}}\approx J^{*}$. This idea constitutes the essence of a broad class of suboptimal schemes.
A popular method of this kind is model predictive control (MPC), which defines a policy through online computation similar to (3). At state $x_{k}$, the MPC method considers the state sequence $x_{k}\dots,x_{k+\ell}$ generated under the control sequence $u_{k},\dots,u_{k+\ell-1}$, where the number $\ell$ is referred to as (the length of) the lookahead steps. The desirability of different sequences are compared using the corresponding sums of the stage costs, plus a function defined on $x_{k+\ell}$. In addition, the state $x_{k+\ell}$ may also be restricted within a set. The function and the constraint associated with $x_{k+\ell}$ are called terminal cost and terminal constraint, respectively. Collectively, they aim to approximate the optimal cost $J^{*}$. The resulting optimization problem with $\ell$-step lookahead can be solved efficiently online. Suppose the minimum of the problem is attained at $\tilde{u}_{k},\dots,\tilde{u}_{k+\ell-1}$, MPC policy $\tilde{\mu}$ is defined by setting $\tilde{\mu}(x_{k})=\tilde{u}_{k}$. When $\ell$ is sufficiently large, and the terminal cost and constraint have certain properties, the resulting MPC policy can be close to optimal.
@@
Taking inspiration from parallel rollout, we propose a new MPC scheme, which we call parallel MPC. It requires solving multiple MPC lookahead minimizations at each time, where each minimization can involve different lookahead steps, and terminal cost and/or constraint. Compared with the MPC scheme noted earlier, the policy $\tilde{\mu}$ defined by parallel MPC has better performance guarantees. Let us use the following scalar example to illustrate the scheme.
@@
(Parallel MPC for a Scalar Problem) Let $X=\Re$, $U(x)=U=[-1,1]$ for all $x$, and $f(x,u)=2x+u$. Assume that $g(x,u)=x^{2}+u^{2}$ if $|x|\leq 2$, and $g(x,u)=\infty$ otherwise. For this problem, an example of parallel MPC can be described as follows: at state $x_{k}$, solve the following two optimization problems
@@
where the constraints corresponding to the dynamics are omitted. Then the optimal values and the controls $\tilde{u}_{k}$ that attain the minimum in the two problems are computed, and the control that corresponds to the minimum of the two optimal values is applied by parallel MPC.
@@
    We show that parallel MPC has a better performance guarantee than the MPC method involving a single lookahead minimization.
@@
    We establish the validity of parallel implementations of the method, which harness the power of multiple computing units.
@@
In Section 2 we provide background and references, which place in context our results in relation to the literature. In Section 3, we describe our MPC algorithm and its variants, and we provide analysis. In Section 4, we demonstrate the validity of our algorithm through computational examples.
@@ 2 Background
The study of infinite horizon optimal control with deterministic dynamics and nonnegative cost has a long and rich history. Exact solutions have been derived for some special types of problems, see, e.g., [2, 21, 26]. For a rigorous analysis that applies more generally, see [6], where the properties of $J^{*}$, as well as algorithmic behaviors for exact solution methods, are addressed.
Despite these extensive research efforts, many deterministic optimal control problems remain intractable. As a result, one often resorts to approximate solution methods. For problems with continuous state and control spaces under constraints, MPC has had abundant applications with wide success. The central themes of MPC analysis are maintaining the feasibility of the $\ell$-step lookahead minimization problems upon reaching new states, and the stability of the resulting closed-loop system under the obtained policy. The former question is typically addressed by introducing a terminal constraint with an invariance property [3]. The terminal cost is often designed to fulfill a condition, known as the Lyapunov inequality, to provide the stability guarantees [29]. Since the early work [22], there have been many MPC variants with these desired properties. Among them, examples involving multiple terminal constraints and/or costs include [28, 16, 24]. Yet these methods lack of generality or flexibility compared with parallel MPC introduced here. Our method is also different from the MPC schemes reported in [18, 19], which focus on solving efficiently a single lookahead minimization via parallel computation. In contrast, our methods involves multiple lookahead minimization problems solved in parallel, and our goal is to improve the performance of the MPC policy measured by its cost function.
@@
Although MPC and parallel rollout are proposed in different contexts, these methods can be unified through a conceptual framework that is based on DP; see the monograph [10] and the surveys [11, 12]. In particular, it is shown that both MPC and parallel rollout are one step of the Newton’s method applied to compute the optimal cost function $J^{*}$, which is responsible for the wide success of these methods. Moreover, the analysis of MPC and rollout can be unified by using a condition that connects to the Lyapunov inequality, as is done in [7, Sections 6.1 and 6.5]. We will use this framework to introduce and analyse the parallel MPC method.
@@
    For all stationary policies $\mu$, if a nonnegative function $\bar{J}:X\to[0,\infty]$ satisfies
    $g\big{(}x,\mu(x)\big{)}+\bar{J}\big{(}f\big{(}x,\mu(x)\big{)}\big{)}\leq\bar{J}(x),\,\forall x\in X,$
    then $J_{\mu}(x)\leq\bar{J}(x)$ for all $x\in X$.
@@ 3 Main results
Let us formally state our parallel MPC algorithm and analyze its properties. We impose the following assumption, which will allow us to replace $\inf$ by $\min$ thereafter. The assumption remains valid throughout this section and thus is omitted in the subsequent theoretical statements.
@@
As long as Assumption 3.1 holds, our method and analysis in this section will be valid. This is true regardless of the nature of the state and control spaces. On the other hand, different types of state and control spaces, being continuous or discrete, may impact the validity of the assumption, and can determine the ways in which we solve the lookahead minimization problems involved in our method.
@@ 3.1 Description of Parallel MPC
Consider the optimal control problem with dynamics (1) and stage cost (2). We assume that we have access to $p$ computational units as well as a central unit. Given the current state $x_{k}$, computational unit $i=1,\dots,p$ solves the problem
@@
where $\ell_{i}$ is some positive integer, which can vary with $i$, and $J_{i}\in\mathcal{L}(X)$ is a function computed off-line in explicit form. The optimal value of the problem obtained by the $i$th unit is denoted by $\tilde{J}_{i}(x_{k})$, and the corresponding minimizing sequence is denoted by $(\tilde{u}^{i}_{k},\tilde{u}^{i}_{k+1},\dots,\tilde{u}^{i}_{k+\ell_{i}-1})$. The values $\tilde{J}_{i}(x_{k})$ and $\tilde{u}_{k}^{i}$ are then sent to the central unit that computes
@@
(Parallel MPC for a Scalar Problem Revisited) In the parallel MPC discussed in Example 1.1, there are $p=2$ computing units. The lookahead minimizations solved in the two units are:
@@
where $\ell_{1}=2$, and $\ell_{2}=1$. The function $J_{1}(x)$ takes the value 00 if $x=0$ and infinity otherwise. Similarly, $J_{2}(x)=(13/3)x^{2}$ if $|x|\leq 2/3$ and infinity otherwise.
@@
In what follows, we will discuss two implementation approaches of parallel MPC, analyse the properties of the corresponding policy $\tilde{\mu}$, and provide a variant of the scheme. For the method defined by (7) and (8), we will use $\ell$ to denote the minimum value of the lookahead steps, i.e., $\ell=\min_{i}\{\ell_{i}\}$. In addition, we define $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$. As a result, the minimizations (7) involving different lookahead steps can be transformed to the ones that have a common lookahead step $\ell$, where the functions used to approximate $J^{*}$ are $\bar{J}_{i}$. This transformation greatly simplifies our analysis, and facilitates the understanding of our scheme.
@@ 3.2 Justification of Parallel MPC
Then Assumption 3.1 asserts that these sets are always nonempty. In addition let $\{\bar{J}_{i}\}_{i=1}^{p}\subset\mathcal{E}^{+}(X)$ be a finite collection of nonnegative functions, and $\bar{J}$ be their pointwise minimum
@@
    $\tilde{U}(T^{\ell-1}\bar{J},x)=\cup_{i\in\tilde{I}(\ell,x)}\tilde{U}(T^{\ell-1}\bar{J}_{i},x)$
@@
  • (a)
    Suppose $\ell=1$. For every fixed $i$, by (11) we have that $\bar{J}(x)\leq\bar{J}_{i}(x)$ for all $x$. By the monotonicity property (6), using $\bar{J}$ and $\bar{J}_{i}$ in place of $J$ and $J^{\prime}$, and taking minimum on both sides, we have
@@
    For the converse, we note that for every $x$, $\tilde{U}(\bar{J},x)$ is nonempty by Assumption 3.1. Let $\tilde{u}$ be an arbitrary element of $\tilde{U}(\bar{J},x)$. Then
@@
  • (b)
    We show that the result holds for $\ell=1$. The induction part is similar to that of the proof for part (a) and is thus omitted. Given an $x$, let $\tilde{u}\in\tilde{U}(\bar{J},x)$, then we have that
@@
    which gives the desired result.
@@
In view of Prop. 3.1(a), we can interpret the proposed scheme (9) as solving $(T^{\ell}\bar{J})(x_{k})$ for some integer $\ell$ and some function $\bar{J}$. To see this, define $\ell=\min_{i}\ell_{i}$, and $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$. Moreover, Prop. 3.1(b) establishes that the unit $\tilde{i}$ selected by the central unit with the smallest value $\tilde{J}_{i}(x_{k})$ must be an element of $\tilde{I}(\ell,x_{k})$. In addition, the first control $\tilde{u}_{k}^{\tilde{i}}$ must be an element of $\tilde{U}(T^{\ell-1}\bar{J},x_{k})$. Therefore, the proposed algorithm solves the problem $(T^{\ell}\bar{J})(x_{k})$ with $\ell$ and $\bar{J}$ defined accordingly. As a result of Prop. 3.1, analysis on the parallel MPC can be transformed to studying $T^{\ell}\bar{J}$. In what follows, we will show that $\bar{J}$ inherits the desired properties of $J_{i}$, thus resulting in the desired behavior of the policy $\tilde{\mu}$.
@@
Let $\{J_{i}\}_{i=1}^{p}\subset\mathcal{L}(X)$ where $p$ is a positive integer, and $\{\ell_{i}\}_{i=1}^{p}$ be a collection of $p$ positive integers. In addition, let $\ell=\min_{i}\ell_{i}$, $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$, and $\bar{J}$ be defined (11). Then $\bar{J}\in\mathcal{L}(X)$.
@@
where the last inequality follows as $\bar{J}\in\mathcal{L}(X)$, by Prop. 3.2. Applying Prop. 2.1(b) with $\bar{J}$ as $T^{\ell-1}\bar{J}$ here, we have $J_{\tilde{\mu}}(x)\leq T^{\ell-1}\bar{J}(x)$. By the monotonicity property (6), we have
@@
The proposition states that the cost function of the parallel MPC policy $\tilde{\mu}$ is upper bounded by the minimum of the values $\tilde{J}_{i}$ computed via individual lookahead. Our computational study, reported in Section 4, also shows that the policy $\tilde{\mu}$ outperforms the MPC policies defined by the individual lookahead minimization.

### v3

@@ Abstract
In this note, we consider infinite horizon optimal control problems with deterministic systems. Since exact solutions to these problems are often intractable, we propose a parallel model predictive control (MPC) method that provides an approximate solution. Our method computes multiple lookahead minimization problems at each time, where each minimization may involve a different number of lookahead steps, and terminal cost and constraint. The policy computed via parallel MPC applies the first control of the lookahead minimization with the lowest cost. We show that the proposed method can harnesses the power of multiple computing units. Moreover, we prove that the policy computed via parallel MPC has better performance guarantee than that computed via the single lookahead minimization involved in parallel MPC.
@@ 1 Introduction
$\mu^{*}(x)\in\arg\min_{u\in U(x)}\big{\{}g(x,u)+J^{*}\big{(}f(x,u)\big{)}\big{\}},\quad\forall x\in X,$ (3)
assuming the minimum in (3) can be attained for all $x$.
In the context of dynamic programming (DP), the goal is to obtain an optimal policy $\mu^{*}$ and optimal cost function $J^{*}$. However, computing $\mu^{*}$ and $J^{*}$ is often intractable. To obtain an approximate solution, one may use some known function $J$ in place of $J^{*}$ and compute a policy $\tilde{\mu}$ via a minimization calculation similar to (3) at each time. If $J$ resembles $J^{*}$ in some way, we may expect the obtained policy $\tilde{\mu}$ close to optimal, in the sense that $J_{\tilde{\mu}}\approx J^{*}$. This idea constitutes the essence of a broad class of suboptimal schemes.
A popular method of this kind is model predictive control (MPC), which defines a policy through online computation similar to (3). [^1] We refer to computations as “online” if they are executed in real time, and as “off-line” if they are performed in advance of online operations. At state $x_{k}$, the MPC method considers the state sequence $x_{k}\dots,x_{k+\ell}$ generated under the control sequence $u_{k},\dots,u_{k+\ell-1}$, where the number $\ell$ is referred to as (the length of) the lookahead steps. The desirability of different sequences are compared using the corresponding sums of the stage costs, plus a function defined on $x_{k+\ell}$. In addition, the state $x_{k+\ell}$ may also be restricted within a set. The function and the constraint associated with $x_{k+\ell}$ are called terminal cost and terminal constraint, respectively. Collectively, they aim to approximate the optimal cost $J^{*}$. The resulting optimization problem is simpler than computing $J^{*}$, and may be possible to solve online. Suppose the minimum of the problem is attained at $\tilde{u}_{k},\dots,\tilde{u}_{k+\ell-1}$, MPC policy $\tilde{\mu}$ is defined by setting $\tilde{\mu}(x_{k})=\tilde{u}_{k}$. When $\ell$ is sufficiently large, and the terminal cost and constraint have certain properties, the resulting MPC policy can be close to optimal.
@@
Taking inspiration from parallel rollout, we propose a new MPC scheme, which we call parallel MPC. It requires solving multiple MPC lookahead minimizations at each time, where each minimization can involve different lookahead steps, and terminal cost and/or constraint. Compared with the MPC scheme noted earlier, the policy $\tilde{\mu}$ defined by parallel MPC has better performance guarantees. Let us consider the following scalar example to illustrate the scheme.
@@
(Parallel MPC for a Scalar Problem) Let $X=\Re$, $U(x)=U=[-1,1]$ for all $x$, and $f(x,u)=2x+u$. Assume that $g(x,u)=x^{2}+u^{2}$ if $|x|\leq 2$, and $g(x,u)=\infty$ otherwise. For this problem, an example of parallel MPC can be described as follows: at state $x_{k}$, solve the following two optimization problems:
@@
with additional constraints $x_{k+1}=2x_{k}+u_{k}$, and $x_{k+2}=2x_{k+1}+u_{k+1}$; cf. Eq. (1). The procedure through which the terminal costs and constraints are computed will be explained in Example 3.3. Then the optimal values and the controls $\tilde{u}_{k}$ that attain the minimum in the two problems are computed, and the control that corresponds to the minimum of the two optimal values is applied by parallel MPC.
@@
    We show that parallel MPC can harnesses the power of multiple computing units.
@@
    We prove that the policy defined by parallel MPC has a better performance guarantee than that defined by the single lookahead minimization involved in parallel MPC.
@@
The paper is organized as follows. In Section 2 we provide background and references, which place in context our results in relation to the literature. In Section 3, we describe our MPC algorithm and its variants, and we provide analysis. In Section 4, we demonstrate our method through computational examples.
@@ 2 Background
The study of infinite horizon optimal control with deterministic systems and nonnegative cost has a long and rich history. Exact solutions have been derived for some special types of problems, see, e.g., [2, 21, 27, 24]. For a rigorous analysis that applies generally, see [6], where the properties of $J^{*}$, as well as algorithmic behaviors for exact solution methods, are addressed.
Despite these extensive research efforts, many deterministic optimal control problems remain intractable. As a result, one often resorts to approximate solution methods. For problems with continuous state and control spaces under constraints, MPC has had abundant applications with wide success. The central themes of MPC analysis are maintaining the feasibility of the $\ell$-step lookahead minimization problems upon reaching new states, and the stability of the resulting closed-loop system under the obtained policy. The former question is typically addressed by introducing a terminal constraint with an invariance property [3]. The terminal cost is often designed to fulfill a condition, known as the Lyapunov inequality, to provide the stability guarantees [30]. Since the early work [22], there have been many MPC variants with these desired properties. Among them, examples involving multiple terminal constraints and/or costs include [29, 16, 25]. Yet these methods lack generality or flexibility compared with parallel MPC introduced here. Our method is also different from the MPC schemes reported in [18, 19], which focus on solving efficiently a single lookahead minimization via parallel computation. In contrast, our methods involves multiple lookahead minimization problems solved in parallel, and our goal is to improve the performance of the MPC policy measured by its cost function.
@@
Although MPC and parallel rollout are proposed in different contexts, these methods can be unified through a conceptual framework that is based on DP; see the monograph [10] and the surveys [11, 12]. In particular, it is shown that both MPC and parallel rollout are one step of the Newton’s method applied to compute the optimal cost function $J^{*}$, which is responsible for the wide success of these methods. Moreover, the analysis of MPC and rollout can be unified by using a condition that connects to the Lyapunov inequality, as is done in [7, Sections 6.1 and 6.5]. We will use this framework to introduce and to analyze the parallel MPC method.
@@
    For all stationary policies $\mu$, if a nonnegative function $J:X\to[0,\infty]$ satisfies
    $g\big{(}x,\mu(x)\big{)}+J\big{(}f\big{(}x,\mu(x)\big{)}\big{)}\leq J(x),\ \ \forall x\in X,$
    then $J_{\mu}(x)\leq J(x)$ for all $x\in X$.
@@ 3 Main Results
Let us formally state our parallel MPC algorithm and analyze its properties. We impose the following assumption, which will allow us to replace $\inf$ by $\min$ thereafter. The assumption remains valid throughout this work and thus is omitted in the subsequent theoretical statements.
@@
Note that our method and analysis is valid under Assumption 3.1, regardless of the nature of the state and control spaces. On the other hand, different types of state and control spaces, being continuous or discrete, may impact the validity of the assumption, and can determine the ways in which we solve the lookahead minimization problems involved in our method.
@@ 3.1 Description of Parallel MPC
Consider the optimal control problem with state equation (1) and stage cost (2). We assume that we have access to $p$ computational units as well as a central unit. Given the current state $x_{k}$, computational unit $i=1,\dots,p$ solves the problem
@@
where $\ell_{i}$ is some positive integer, which can vary with $i$, and $J_{i}\in\mathcal{L}(X)$ is a function computed off-line in explicit form. [^2] We have restricted our attention to the cases where the functions $J_{i}$ are known in explicit forms. Regarding the situation where the functions $J_{i}$ can only be evaluated online through simulation, as in parallel rollout given in [8, Section 5.2], our method and analysis to be introduced shortly applies in principle. However, in such a case, the lookahead minimization involved in our method can be solved only if the control space is discrete. The optimal value of the problem obtained by the $i$th unit is denoted by $\tilde{J}_{i}(x_{k})$, and the corresponding minimizing sequence is denoted by $(\tilde{u}^{i}_{k},\tilde{u}^{i}_{k+1},\dots,\tilde{u}^{i}_{k+\ell_{i}-1})$. The values $\tilde{J}_{i}(x_{k})$ and $\tilde{u}_{k}^{i}$ are then sent to the central unit that computes
@@
In the parallel MPC discussed in Example 1.1, there are $2$ computing units. The lookahead minimization problems solved in the two units are:
@@
with additional constraints $x_{k+1}=2x_{k}+u_{k}$, and $x_{k+2}=2x_{k+1}+u_{k+1}$. Here $\ell_{1}=2$, and $\ell_{2}=1$. The function $J_{1}(x)$ takes the value 00 if $x=0$ and infinity otherwise. Similarly, $J_{2}(x)=(13/3)x^{2}$ if $|x|\leq 2/3$ and infinity otherwise.
@@
In what follows, we will show the equivalence between parallel MPC and an MPC scheme involving a single lookahead minimization with a suitably defined terminal cost. We will also analyze the properties of the corresponding policy $\tilde{\mu}$, and provide a variant of the scheme. For parallel MPC defined by (7) and (8), we denote by $\ell$ the minimum value of the lookahead steps, i.e., $\ell=\min_{i}\{\ell_{i}\}$. In addition, we define $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$. As a result, the minimization (7) involving different lookahead steps can be transformed to the ones that have a common lookahead step $\ell$ with terminal costs $\bar{J}_{i}$.
@@ 3.2 Justification of Parallel MPC
Assumption 3.1 asserts that these sets are always nonempty. In addition let $\{\bar{J}_{i}\}_{i=1}^{p}\subset\mathcal{E}^{+}(X)$ be a finite collection of nonnegative functions, and $\bar{J}$ be their pointwise minimum
@@
    $\tilde{U}(T^{\ell-1}\bar{J},x)=\cup_{i\in\tilde{I}(\ell,x)}\tilde{U}(T^{\ell-1}\bar{J}_{i},x)$ (15)
@@
(a) Suppose $\ell=1$. For every fixed $i$, by (12) we have that $\bar{J}(x)\leq\bar{J}_{i}(x)$ for all $x$. By the monotonicity property (6), using $\bar{J}$ and $\bar{J}_{i}$ in place of $J$ and $J^{\prime}$, and taking minimum on both sides, we have
@@
For the converse, we note that for every $x$, $\tilde{U}(\bar{J},x)$ defined in Eq. (11) is nonempty by Assumption 3.1. Let $\tilde{u}$ be an arbitrary element of $\tilde{U}(\bar{J},x)$. Then
@@
(b) We show that the result holds for $\ell=1$. The induction part is similar to that of the proof for part (a) and is thus omitted. Given an $x$, let $\tilde{u}\in\tilde{U}(\bar{J},x)$, then we have that
@@
which gives the desired result. ∎
@@
In view of Prop. 3.1(a), we can interpret the proposed scheme (9) as solving $(T^{\ell}\bar{J})(x_{k})$ for some integer $\ell$ and some function $\bar{J}$. To see this, define $\ell=\min_{i}\{\ell_{i}\}$, and $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$. Moreover, Prop. 3.1(b) establishes that the unit $\tilde{i}$ selected by the central unit with the smallest value $\tilde{J}_{i}(x_{k})$ must be an element of $\tilde{I}(\ell,x_{k})$. In addition, the first control $\tilde{u}_{k}^{\tilde{i}}$ must be an element of $\tilde{U}(T^{\ell-1}\bar{J},x_{k})$. Therefore, the proposed algorithm solves the problem $(T^{\ell}\bar{J})(x_{k})$ with $\ell$ and $\bar{J}$ defined accordingly. As a result of Prop. 3.1, analysis on the parallel MPC can be transformed to studying $T^{\ell}\bar{J}$. Indeed, the left side of Eq. (14) serves as an analytical device for our investigation on the properties of parallel MPC, while the right side of Eq. (14) represents the actual computational approach we employ. This is because the optimization problem defined by $(T^{\ell}\bar{J})(x)$ is often more challenging to solve compared with that defined by the right side of Eq. (14); see Examples 3.3 and 4.1. Yet, it is easier to analyze the properties of the policy defined by $(T^{\ell}\bar{J})(x)$.
In what follows, we will show that $\bar{J}$ inherits the desired properties of $J_{i}$, thus resulting in the desired behavior of the policy $\tilde{\mu}$.
@@
Let $\{J_{i}\}_{i=1}^{p}\subset\mathcal{L}(X)$ where $p$ is a positive integer, and $\{\ell_{i}\}_{i=1}^{p}$ be a collection of $p$ positive integers. In addition, let $\ell=\min_{i}\{\ell_{i}\}$, $\bar{J}_{i}=T^{\ell_{i}-\ell}J_{i}$, and $\bar{J}$ be defined (12). Then $\bar{J}\in\mathcal{L}(X)$.
@@
where the last inequality follows as $\bar{J}\in\mathcal{L}(X)$, by Prop. 3.2. Applying Prop. 2.1(b) with $T^{\ell-1}\bar{J}$ as $J$, we have $J_{\tilde{\mu}}(x)\leq T^{\ell-1}\bar{J}(x)$. By the monotonicity property (6), we have
@@
Let us denote by $\tilde{\mu}_{i}$ the MPC policy defined by $T^{\ell_{i}}J_{i}$; i.e., the MPC policy computed by the $i$th unit. Classical MPC theory states that $J_{\tilde{\mu}_{i}}(x)\leq\tilde{J}_{i}(x)\leq\infty$ for all $x\in X$. The proposition states that the cost function of the parallel MPC policy $\tilde{\mu}$ is upper bounded by the pointwise minimum of the functions $\tilde{J}_{i}$ computed via individual lookahead. Therefore, $J_{\tilde{\mu}}(x)<\infty$ as long as $\tilde{J}_{i}(x)<\infty$ for some $i$. This implies that the set of the states $x$ where $J_{\tilde{\mu}}(x)<\infty$ is the union of the sets $\{x\in X\,|\,J_{\tilde{\mu}_{i}}(x)<\infty\}$, $i=1,2,\dots,p$.
Remark 3.3.
Under certain conditions, we can show that the cost function $J_{\tilde{\mu}}$ satisfies $J_{\tilde{\mu}}\leq J_{\tilde{\mu}_{i}}$ for some $i$. In particular, if $J_{j}=J_{\tilde{\mu}_{i}}$ for some $j\neq i$, then $\tilde{J}_{j}=T^{\ell_{j}}J_{j}\leq J_{j}=J_{\tilde{\mu}_{i}}$. Such a case arises when the function $J_{j}$ is constructed via collecting state-and-control trajectory data under $\tilde{\mu}_{i}$, as is proposed in [31]. Our computational study, reported in Section 4, also provides empirical evidence that the policy $\tilde{\mu}$ often outperforms the MPC policies defined by the individual lookahead minimization even without such conditions.
@@
We demonstrate these propositions using our running example.
Example 3.2.
Consider the parallel MPC applied to the problem studied in Examples 1.1 and 3.1. Prop. 3.1 states that the minimum of the optimal values computed by the two units, as shown in Example 3.1, equals to the minimum $\tilde{J}(x_{k})$ of the following problem:
$\min_{u_{k}\in U}$ $\quad g(x_{k},u_{k})+\bar{J}(x_{k+1})$ (17)
$\mathrm{s.\,t.}$ $\quad\bar{J}(x_{k+1})=\min\big{\{}\bar{J}_{1}(x_{k+1}),\bar{J}_{2}(x_{k+1})\big{\}},$
$\quad\bar{J}_{1}(x_{k+1})=\min_{\begin{subarray}{c}u_{k+1}\in U\\ 2x_{k+1}+u_{k+1}=0\end{subarray}}g(x_{k+1},u_{k+1}),$
$\quad\bar{J}_{2}(x_{k+1})=\begin{cases}13/3x_{k+1}^{2},&\text{if }|x_{k+1}|\leq 2/3,\\ \infty,&\text{otherwise;}\end{cases}$
cf. Eq. (14). Moreover, the control $\tilde{u}_{k}$ that attains the minimum in problem (17) also attains the minimum of the problem computed by the unit with smaller minimum values in problems (10); cf. Eq. (15). Clearly, problems (10) are easier to solve than problem (17).
Taking as given that $\bar{J}_{1},\bar{J}_{2}\in\mathcal{L}(X)$ (we will show this later), Props. 3.2 and 3.3 state that the policy $\tilde{\mu}$ defined by $\tilde{\mu}(x_{k})=\tilde{u}_{k}$ satisfies $J_{\tilde{\mu}}(x_{k})\leq\tilde{J}(x_{k})$.